The sides of a triangle are in the ratio of 2:4:5. If the perimeter of the triangle is 88 meters, determine the length of the each side.
Answer by Alan3354(69443)

2+4+5 = 11
88/11 = 8
----------
2*8 = 16
4*8 = 32
5*8 = 40